Kagura:

Determine the equivalent system for the given system of equations. (1 point) 4x − 5y = 2 10x − 21y = 10 4x − 5y = 2 3x − y = 4 4x − 5y = 2 24x − 47y = 22 4x − 5y = 2 10x + 3y = 15 4x − 5y = 2 14x + 26y = 12

Kagura:

hallo

luhivqqcherry:

4x−5y=2 24−47=22 24x−47y=22 .

Kagura:

b?

luhivqqcherry:

yawpp .

Markthegoat1117:

😂

luhivqqcherry:

im getting mathelete before mark 😠.

Kagura:

oof

luhivqqcherry:

mUhahaha .

luhivqqcherry:

😠.

luhivqqcherry:

😠.

Kagura:

iont even have 50 medals

Kagura:

for math

Markthegoat1117:

How old are you

luhivqqcherry:

@kagura wrote:
iont even have 50 medals
i have like 150 sum medals in math .

luhivqqcherry:

@markthegoat1117 wrote:
How old are you
8yrs old 😠.

Markthegoat1117:

Nah you like 20

luhivqqcherry:

...

Markthegoat1117:

Have to be

luhivqqcherry:

😠.

luhivqqcherry:

stop getting off topic .

Markthegoat1117:

You outta schoo

luhivqqcherry:

stop getting off topic .

luhivqqcherry:

😠.

Markthegoat1117:

You faster bc you older and not in schoo...im literally sitting in class rn🤣🤣